OpenStudy (anonymous):

A triangle has an area of 10cm^2. The height of the triangle is 4cm longer than three times the base. (((((((Find the height and base of the triangle)))))))

OpenStudy (anonymous):

Area of a triangle= 1/2 base x height 10= 1/2 (b)(3b+4)

OpenStudy (anonymous):

solve that out for b to get your base and then the height is 3b+4

OpenStudy (anonymous):

.5base x height???

OpenStudy (anonymous):

do I want to move everything over to the left except the base?

OpenStudy (anonymous):

isolate the b on one side, so yes

OpenStudy (anonymous):

10= 1/2 (b)(3b+4) 10= 1/2 (3b^2+4b) 20= 3b^2+4b 3b^2+4b-20= 0

OpenStudy (anonymous):

(3b+10)(b-2)= 0 3b=-10 or b=2 b=-10/3

OpenStudy (anonymous):

However, if you think about it logically, can a base ever be negative?

OpenStudy (anonymous):

No it can not

OpenStudy (anonymous):

Thus, the base is 2 and the height is 3(2)+4= 6+4=10

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

the height is 10??

OpenStudy (anonymous):

yes, because the height= 3b+4 and the base is 2 so 3(2)+4 6+4 10

OpenStudy (anonymous):

0r\[-\frac{ 10 }{ 3 }\]

OpenStudy (anonymous):

A base or height can never be negative though

OpenStudy (anonymous):

ok I got it.

OpenStudy (anonymous):

okay cool :)